Suppose a sample of size 12 is taken, resulting in a sample mean of 300 and a sample standard deviation of 30. Determine the Point Estimate and the Standard Error for a 99% Confidence Interval for this sample. Round all calculations to two (2) decimal places.
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 300

sample standard deviation = s = 30

sample size = n = 12

Standard error =  (s /n) = ( 30 / 12) = 8.66